OpenStudy (anonymous):

In a number comprised of two digits, the the tens digit exceeds in 5 the unit digit. If we invert the order of the two digits, we have a new number. This new number added to the previous result in 121. ¿what is the number? Please help

OpenStudy (anonymous):

Let x = units digit y = tens digit y = x + 5 x + 10y + y + 10x = 121 x + 10(x + 5) + x + 5 + 10x = 121 x + 10 x + 50 + x + 5 + 10x = 121 22x + 55 = 121 22x = 66 x = 3 y = 8 83

OpenStudy (anonymous):

Thanks again

OpenStudy (anonymous):

10 point for you